Post this question to forum Bryce runs a 100 meter race. 5 seconds after the race started Bryce is 45 meters from the starting line and reaches his max speed, he runs at this max speed for the rest of the race. Bryce notices that he is 77 meters from the starting line 9 seconds after the race started. a What is Bryces max speed? meters per second Preview b. Suppose Bryce runs for an additional z seconds after reaching his max speed.. i How far will Bryce travel during those additional z seconds? (Preview; ii What is Bryces distance from the starting line 5+ z seconds after the race started? meters # meters Preview c. What is Bryces distance from the starting line z seconds after the race started (provided z 2 5)? meters Preview Submit Question 4. Points possible: 4 Unlimited attempts Score on last attempt 1. Score in gradebook: 1 Message instructor about this question & Ucense

Bruce runs a 100m race. 5 seconds after the race started Bruce is 45 mi from the starting line and reaches his max speed, he runs at his max speed for the rest of the race. Bryce notices that he is 77 meters from the starting line 9 seconds after the race started.

a. what is Bryce's max speed

b. suppose Bryce runs for an additional Z seconds after reaching his max speed

i. how far will Bryce travel during those additional Z seconds

ii. what is Bryce's distance from The Starting Line 5 + z seconds after the race started

c. what is Bryce's distance from The Starting Line X seconds after the race started?
Answer #1

Solution

a) Bryce maximum speed -

Speed = Distance / Time

Distance covered with maximum speed = 77 – 45

= 32 meters

Time taken in maximum speed = 9 – 5

= 4 seconds

Therefore speed = distance / time

= 32/4

                                   = 8 meters/ seconds

b) Bryce runs an additional z seconds after reaching maximum speed.

So his total time = 5+z seconds

i) How far will Bryce travel in during those additional z seconds

DistanceD = Speed S * Time T

Speed S = 8 m/s

Time T = z seconds

Therefore D =8 * z

                = 8z meters

ii) Bryce distance from the starting line 5+z seconds –

Distance D = speed S* time T

Speed = 8 m/s

Time = 5+z seconds

Distance = 8*(5+z)

                = 45+8z meters

c) Bryce distance from the starting line x seconds after the race started ( provided x>_ 5)

Distance D = Speed S* Time T

Speed = 8 m/s

Time = x seconds

D = S*T

= 8 * x

= 8x meters

Please note : If we consider the time as 5+x seconds then the calculation will be

D = 8*(5+x)

= 45+8x meters